# ECC Analyzer: Extract Trading Signal from Earnings Conference Calls using Large Language Model for Stock Performance Prediction

### v1

@@ D.1 Prompt for Summarizing Earnings Conference Call Segments
D.2 Prompt for Creating an Overview Summary from Earnings Conference Call Segments
@@
D.3 Prompt for Extracting Focus And Explore Its Impact
  • Listen to the Call
@@ D.3 Prompt for Extracting Focus And Explore Its Impact
    Identify statements or discussions that involve significant financial metrics, strategic initiatives, new products or markets, regulatory impacts, or any notable shifts in operations. These are potential focus points that could influence investor perceptions and stock price.
@@
    Pre and Post Analysis: Examine stock price movements immediately before and after the call to capture initial reactions.
@@
D.4 Prompt for Analyzing Sentiment of Focus Points and Supporting Evidence from an Earnings Conference Call
  • Analyze Sentiment for Each Focus Point
    Apply the sentiment analysis to the text surrounding each focus point. Pay attention to the language used, such as positive, negative, or neutral descriptors, and the intensity of the language. Please assign a continuous decimal sentiment score to each event, ranging from -1 to 1. A score of -1 represents a highly negative sentiment, 0 indicates neutrality, and 1 signifies a highly positive sentiment.
  • Context Consideration
    Consider the context in which each point was discussed. Assess whether the sentiment is directly related to the focus point or influenced by broader discussion themes.
  • Document Supporting Quotes
    For each focus point, document specific quotes or statements from the call that illustrate the sentiment. Note the speaker and their role to add credibility to the sentiment analysis.
D.5 Prompt for Analyzing Audio Features of Focus Points from an Earnings Conference Call
  • Tone and Pitch Analysis
    Examine variations in tone and pitch within each audio segment. Look for patterns that might indicate emphasis, uncertainty, confidence, or stress.
  • Volume and Speech Rate
    Measure changes in volume and variations in speech rate. High volume and rapid speech may indicate areas of strong emotion or importance.
  • Pause Patterns
    Identify the frequency and duration of pauses, which can provide insights into the speaker’s thought process or hesitation.

### v2

@@ A.1.1. Prompt for Summarizing Earnings Conference Call Segments
A.1.2. Prompt for Creating an Overview Summary from Earnings Conference Call Segments
@@
A.1.3. Prompt for Extracting Important Sentences.
  • Check to the Call
@@
    Identify statements or discussions that involve significant financial metrics, strategic initiatives, new products or markets, regulatory impacts, or any notable shifts in operations. These are potential focus points that could influence investor perceptions and stock prices.
@@
    Pre and Post-Analysis: Examine stock price movements immediately before and after the call to capture initial reactions.
@@
A.2. Design of Question Bank
Table 3. The designed Question Bank
Focus Category Focus Item Questions
Financial Indicator Dividend Q1: Did this company pay the investors dividend? Q2: Have there been any increases or decreases in the stock dividends? If yes, what is the rate at which the dividends have been increasing? Q3: What type of dividend did the company pay?
Revenue Q1: What was the company’s reported revenue for the past quarter, and how does it compare to the same quarter in the previous year? Q2: What factors influenced the company’s revenue performance this quarter? Q3: What are the company’s revenue forecasts for the upcoming quarters, and what strategies are in place to achieve these targets?
Return Q1: What was the company’s net profit margin for this quarter, and how has it changed from the previous quarter or year? Q2: What is the company’s Return on Equity (ROE) and Return on Assets (ROA) for this period? Q3: How does the company plan to enhance shareholder value in the upcoming periods? Are there any dividends or buybacks planned?
Earnings Q1: Have there been any increases or decreases in the earnings? If yes, what is the rate at which the earnings have been increasing or decreasing? Q2: What is the outlook provided by the executives of this company in relation to the future earnings growth? Q3: Are the earnings above or below compared to the expectations? Are they attractive compared to the peers?
Employee Manager Salary Q1: What percentage of the company’s total expenses is currently allocated to employee salaries, and how has this changed in response to recent business developments? Q2: How does your company’s compensation structure compare with industry standards, particularly in terms of salary, benefits, and bonuses? Q3: What were the average salary increases or decreases across the company this year compared to last year?
Pension Q1: What is the current status of the company’s pension fund, and what were the major changes to its funding status over the past year? Q2: How does the company manage its pension liabilities, and what strategies are in place to address any underfunded positions? Q3: What are the expected impacts of current pension commitments on the company’s future financial performance?
Management Change Q1: Have there been any recent changes in the company’s key management positions, including the CEO, etc. Q2: What were the reasons behind any recent management changes, particularly in the CEO position? Q3: What impact are the recent management changes expected to have on the company’s strategy in the near term?
Cost Operating Costs Q1: What were the total operating costs this quarter compared to the previous quarter? Q2: Which factors contributed to any significant changes in operating costs? Q3: How are you managing operating costs in light of current economic conditions?
Cost of Goods Sold Q1: How did the Cost of Goods Sold change this quarter, and what were the driving factors behind these changes? Q2: What percentage of revenue does the Cost of Goods Sold represent, and how does this compare to industry norms? Q3: Are there any initiatives in place to reduce Cost of Goods Sold without compromising quality?
Marketing and Sales Costs Q1: How much did the company spend on marketing and sales this quarter? Q2: What specific marketing or sales strategies contributed to these costs? Q3: Are there plans to adjust these strategies in the upcoming quarters based on performance?
Expansion Geographic Expansion Q1: What specific regions or markets is the company expanding into, and what factors influenced these selections? Q2: What are the initial costs associated with the geographic expansion and strategies are in place to support it? Q3: What is the projected timeline for the new regional operations to reach profitability?
Product Line Expansion Q1: What new products is the company introducing, and what consumer or market needs do they aim to address? Q2: How will the introduction of these products impact the company’s production costs and overall financial performance? Q3: Are there any expected synergies between the new products and existing products or services?
Market Segmentation Expansion Q1: Which new customer segments is the company targeting, and what research supports this strategic direction? Q2: What marketing strategies will be employed to reach these new segments, and what are the anticipated costs? Q3: What are the growth expectations for these new market segments over the next few years?
Business Business Q1: What are the key projects or initiatives currently being undertaken by the company, and provide a brief overview of what each project entails? Q2: How has the performance of these projects compared to the previous quarter, and how do they stand in relation to competitors in the same sector? What has been the market’s response to these initiatives? Q3: What are the generated revenues from these projects for the current reporting period, and what potential risks could impact their future performance?
Future Outlook Future Outlook Q1: What are the company’s primary strategic goals for the upcoming year, and what key initiatives are planned to achieve these objectives? Q2: How do these future plans align with current industry trends and market demands? Q3: What are the expected financial impacts of these plans on the company’s performance in the short and long term?